sevans:

If a car changes its velocity from 32 km/hr to 54 km/hr in 8.0 seconds, what is its acceleration?

Shadow:

\[acceleration = \frac{ \Delta v }{ \Delta t } = \frac{ v_{f} - v_{i} }{ time }\] Acceleration is the change in velocity, over time. To simply subtract the final velocity by the initial velocity, then divide it by the time it took for that change to occur.

Shadow:

Delta is the triangle symbol, which just means 'change in'

Shadow:

First though, I would convert the velocities to m/s 1km = 1000 meters

Shadow:

and an hour is 3600 seconds

Shadow:

@sevans Do you understand how to proceed?